OpenStudy (anonymous):

a/5= 6/b is an example of a proportion where a ≠ 0 and b ≠ 0. True or False?

OpenStudy (cwrw238):

true

OpenStudy (anonymous):

Can you explain why?

OpenStudy (cwrw238):

a proportion is the name given to a statement that 2 ratios are equal eg 6/2 = 12/4 or 1/3 = 3/9 if the numbers are a and b in your example are 0 then there is no equality 0/5 is not equal 6/0